Be sure to answer all parts.

Calculate the pOH and pH of the following aqueous solutions at 25

°

C:

(a) 0.0775 M LiOH

     pH =

      pOH =



(b) 0.0441 M Ba(OH)2

     pH =

      pOH =



(c) 0.25 M NaOH

     pH =

      pOH
Answer #1

a)

[OH-] = [LiOH] = 0.0775 M

use:

pOH = -log [OH-]

= -log (7.75*10^-2)

= 1.11

use:

PH = 14 - pOH

= 14 - 1.11

= 12.89

Answer:

pH = 12.89

pOH = 1.11

b)

[OH-] = 2*[Ba(OH)2]

= 2*0.0441 M

= 0.0882 M

use:

pOH = -log [OH-]

= -log (8.82*10^-2)

= 1.05

use:

PH = 14 - pOH

= 14 - 1.05

= 12.95

Answer:

pH = 12.95

pOH = 1.05

c)

[OH-] = [NaOH] = 0.25 M

use:

pOH = -log [OH-]

= -log (0.25)

= 0.602

use:

PH = 14 - pOH

= 14 - 0.6021

= 13.40

Answer:

pH = 13.40

pOH = 0.60